OpenStudy (awkwardpanda):

what are the vertex, focus, and directrix of the parabola with equation y=x^2-10x+33?

OpenStudy (anonymous):

yikes vertex is easiest

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which, in your case, is \(-\frac{-10}{2}=5\)

OpenStudy (anonymous):

second coordinate is what you get when you replace \(x\) by \(5\) i get \(8\) so the vertex is \((5,8)\)

OpenStudy (awkwardpanda):

yeah i got the vertex but what's the focus

OpenStudy (anonymous):

that means you can write it as \[y=(x-5)^2+8\] or \[y-8=(x-5)^2\]

OpenStudy (anonymous):

this looks like \(4p(y-k)=(x-h)^2\) (kinda) where \(4p=1\) so \(p=\frac{1}{4}\)

OpenStudy (anonymous):

that makes the focus \(\frac{1}{4}\) unit above the vertex since the vertex is \((5,8)\) the focus is \((5,8+\frac{1}{4})\) or \[(5,\frac{33}{4})\]

OpenStudy (anonymous):

if (god help me) that is not one of your "answer choices" it could be \[(5,8\tfrac{1}{4})\]

OpenStudy (anonymous):

you got it? the directrix is \(\frac{1}{4}\) units down at \(y=\frac{31}{4}\)

OpenStudy (awkwardpanda):

so. what you're saying is that vertex (5,8) focus (5, 8.25) directrix (y=7.75

OpenStudy (anonymous):

lol i thought i had the bases covered with fractions, and mixed numbers but now decimals yes, that is correct

OpenStudy (awkwardpanda):

xD yeesss. thank you, my savior.

OpenStudy (anonymous):

yw!